# suppReferences

### 17 Mar 2024 v1

@@ Abstract
Misinformation undermines public trust in science and democracy, particularly on social media where inaccuracies can spread rapidly. Experts and laypeople have shown to be effective in correcting misinformation by manually identifying and explaining inaccuracies. Nevertheless, this approach is difficult to scale, a concern as technologies like large language models (LLMs) make misinformation easier to produce. LLMs also have versatile capabilities that could accelerate misinformation correction—however, they struggle due to a lack of recent information, a tendency to produce plausible but false content and references, and limitations in addressing multimodal information. To address these issues, we propose Muse, an LLM augmented with access to and credibility evaluation of up-to-date information. By retrieving contextual evidence and refutations, Muse can provide accurate and trustworthy explanations and references. It also describes visuals and conducts multimodal searches for correcting multimodal misinformation. We recruit fact-checking and journalism experts to evaluate corrections to real social media posts across 13 dimensions, ranging from the factuality of explanation to the relevance of references. The results demonstrate Muse’s ability to correct misinformation promptly after appearing on social media—overall, Muse outperforms GPT-4 by 37% and even high-quality corrections from laypeople by 29%. This work underscores the potential of LLMs to combat real-world misinformation effectively and efficiently.
@@ Introduction
Misinformation—including misinformed, inaccurate, and misleading information—has far-reaching and detrimental effects on individuals and society[1, 2, 3, 4, 5]. It erodes public trust in government, decreases civil engagement in elections, and has been viewed as a threat to democracy[6, 3, 7]. Evidence has demonstrated that election misinformation helped fuel the January 6th attack on the U.S. Capitol, where five people died and more than 100 police officers were injured[8]. Misinformation also drastically increases during outbreaks and disasters, as seen with the “infodemic” during the COVID-19 pandemic, which significantly increased vaccine hesitancy[9, 10]. The result is tragic; an analysis showed that COVID-19 vaccines could have prevented at least 232,000 of COVID-19-associated deaths between May 2021 and September 2022 in the U.S. alone[11]. Concerns over misinformation on social media have been particularly significant[12, 13, 14], as the social media context interferes with truth discernment, where users post content without professional moderation and often consume news in a hasty and distracted way[15, 16].
Fortunately, high-quality and prompt misinformation correction, which broadly includes identifying what part(s) of content is (or is not) misinformation and explaining why that part of the content is misinformed, inaccurate, or misleading, has effectively reduced people’s false beliefs[17, 4, 18]. To this end, domain experts (e.g., on FactCheck.org) have been instrumental in providing accurate, detailed, and well-written explanations of whether and why a claim can be inaccurate with references. Subsequently, other research has shown that groups of laypeople can also identify and produce effective explanations to correct misinformation if their assessments are strategically aggregated[18, 19, 20]. Nevertheless, combating misinformation in the digital and generative artificial intelligence (AI) era cannot solely rely on human intelligence[21, 22]. The number of active fact-checking sites has remained stagnant since 2022[23]. Even though crowdsourcing is more scalable than expert-led activities, keeping pace with millions of posts daily on social media is impossible. As Brandolini’s law indicates[24], correcting misinformation is laborious—often requiring one to search and evaluate the credibility of related articles with evidence completing or refuting the claims in the given piece and write responses as justifications—whereas creating and spreading misinformation is easy. The absence of sufficient coverage markedly diminishes the impact of correction and, as the implied truth effect suggests, may even increase the perceived accuracy of misinformation that escapes correction[25].
@@
While generative AI models such as LLMs (e.g., OpenAI’s GPT series[26]) raise concerns that they facilitate the creation of misinformation, they also potentially make scaling up the correction of misinformation possible. Recent LLMs have exhibited proficiency in generating fluent and coherent text, laying a foundation for producing explanations that the public can understand. Indeed, LLMs have revolutionized the field of AI and presented remarkable capabilities across domains and tasks[27, 28, 29]. However, accurate and trustworthy misinformation correction is inseparable from accessing up-to-date and reliable information, providing accurate references to back up claims, and addressing textual and visual information, all areas where existing LLMs fall short[30]. For instance, GPT-4[26] (as of March 2023) and MisinfoCorrect[31] lack access to timely updated knowledge and are thereby ill-equipped to combat misinformation on emerging topics. They either do not provide or “hallucinate” references, which can be fabricated or irrelevant[30, 32]. A growing body of literature has focused on retrieval-augmented LLMs, which can retrieve up-to-date information from Wikipedia or the entire Internet[33, 34, 35, 36, 37, 38, 39, 40]. However, their retrieval does not explicitly consider the factuality and bias of retrieved sources, posing risks of generating inaccuracies and backfiring (i.e., reinforcing rather than reducing false beliefs)[34, 41, 42]. These LLMs, as well as MisinfoCorrect[31], also struggle with counteracting multimodal misinformation due to their nonacceptance of visual inputs. Finally, understanding the quality of a correction necessitates a comprehensive human evaluation[30]. However, prior uses of LLMs for misinformation correction were evaluated by non-experts on limited aspects of their generated responses (e.g., the general quality of responses[31])[40], which is insufficient to comprehensively understand their performance in identifying and explaining inaccuracies, generated text, and references.
@@
In this article, we propose Muse, a scalable approach for multimodal misinformation correction. Muse makes use of an LLM and augments it with the ability to handle visuals, access timely, factual, and credible knowledge on the web, retrieve evidence that refutes or contextualizes misinformation, and generate clear explanations with accurate and trustworthy references. We recruited experts in fact-checking and journalism to comprehensively evaluate corrections made by Muse, and compare them to baselines including GPT-4 and high-quality corrections written by the collective efforts of laypeople on social media. Our assessment measures the overall quality of corrections, specifically defined as the explicitness, accuracy, comprehensiveness, and informativeness of corrections, the relevance, factuality, fluency, coherence, and toxicity of generated text, and the reachability, relevance, and credibility of references. We find that overall Muse outperforms GPT-4 by 37% and high-quality corrections from laypeople by 29%, demonstrating Muse’s ability to correct both textual and visual misinformation across a broad range of domains, including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine, promptly after appearing on social media.
@@
Muse is designed to automatically generate a correction to a given piece of social media content that is potentially misinformation. As the input, the content can contain text with or without visual information. As the output, the correction should identify what part(s) of the content is (or is not) misinformation, explain why that part of the content is misinformation, and provide links as references. The general pipeline of Muse is illustrated in Fig. 1.
@@
We start by introducing the details of Muse with a piece of text-only misinformation as the input. First, Muse generates queries based on an LLM from the misinformation (Fig. 1b; Methods). Each query acts as the input of a web search engine to access timely updated web content and obtain a list of web links directly relevant to the query (Fig. 1b; Methods). After scraping the content from these web links, Muse calculates their direct relevance to the misinformation and removes irrelevant web pages (Fig. 1b; Methods). Then, Muse determines the credibility of web pages by looking up their publishers’ factuality and bias ratings and selects pages with high factuality and minimal bias (Fig. 1c; Methods). Next, Muse leverages an LLM to extract sentences or paragraphs from each of the web pages as evidence, i.e., the text either explicitly refutes the misinformation or provides additional context (Fig. 1d; Methods). Finally, Muse generates a correction to the misinformation by providing an LLM with the extracted pieces of evidence and their source web links (Fig. 1d; Methods).
@@
Note that misinformation, especially when it comes to social media, often contains extraneous information that is irrelevant to a correction, including unverifiable opinions or emojis, such as the textual content of the misinformation in Fig. 1. Therefore, generating queries instead of simply using the text of the post improves web searches as a way of denoising misinformation; see example queries in Fig. 1b. Meanwhile, generating multiple queries helps decompose a post, which may have multiple claims that each needs to be corrected, whereas generating one query may overlook some misinformed claims (Supplementary Fig. S1). Another concern may arise from filtering retrieved web pages based on how relevant the content is to the misinformation: theoretically, including all retrieved web pages increases the amount of extracted evidence, which may not hurt and perhaps even benefit correction. However, the increase in selected web pages drastically elevates the expense of Muse (Methods). We also observe that retrieved web pages with relatively low relevance can increase the prevalence of hallucinations when generating corrections (Supplementary Fig. S2). Moreover, as illustrated in Fig. 1c, Muse ranks the selected web pages by their publishers’ factuality and bias. It starts extracting evidence from pages with the highest factuality and least bias and then continues down the ranking, stopping when it has obtained sufficient refutations (i.e., at least two web pages were found to refute the misinformation).
@@
When the input is a piece of multimodal misinformation with textual and visual information, Muse first describes each image in text (Fig. 1a) so that the misinformation can be handled by any LLM on downstream tasks, including query generation (Fig. 1b), evidence extraction (Fig. 1d), and correction generation (Fig. 1d). Specifically, Muse augments image captioning models developed to describe an image in natural language with recognizing celebrities and optical characters based on an LLM (Fig. 1a; Methods). Compared to existing image captioning models that capture global features of images, Muse produces more informative descriptions with additional features, including identification of celebrities and embedded text—information crucial for making accurate corrections[43]. For example, even a state-of-the-art image captioning model[44] may describe the visual misinformation in Fig. 1 as simply “list of banned books in Florida”. The description overlooks the listed titles of books that are essential for the visual misinformation’s correction (see more examples in Supplementary Fig. S3). In addition, Muse conducts multimodal search on the web and computes multimodal relevance to filter out irrelevant web pages (Fig. 1b; Methods).
@@
Muse’s evaluation was based on X Community Notes data. Community Notes (communitynotes.twitter.com, formerly Twitter Birdwatch) was launched in 2021 to empower people on X to collaboratively add context to potentially misleading posts. It has been shown to effectively reduce the spread of misinformation[20]. Community Notes operates by enabling people on X, often laypeople, to identify whether a tweet is potentially misleading, write free-response corrections of any tweet, and assess whether other people’s corrections are helpful or unhelpful. The platform strategically aggregates people’s assessments with a transparent algorithm such that a correction has a higher helpfulness score if more people with diverse backgrounds (e.g., with different political ideologies) agree that the correction was helpful[20]. Besides canceling out inaccurate judgments made by some individuals, the algorithm has been shown to empirically address the “tyranny of the majority” and coordinated attempts at gaming[20, 19]. A correction with a sufficiently high helpfulness score is then displayed on the corresponding tweet on X and visible to the public[20] (Methods; Supplementary Fig. S4).
@@
We selected all potentially misleading tweets from Community Notes that have at least two corrections, where one correction has a high helpfulness score and the other has an average helpfulness score (Methods). We further generated corrections to these tweets based on Muse and GPT-4 (as of June 2023) (Discussion; Methods). Experts in fact-checking and journalism were recruited, onboarded, and asked to evaluate the quality of corrections by various approaches to the same tweet (Methods). Experts were blinded to which approach had generated each correction. The evaluation contains 13 specific criteria, covering how well the correction identifies and explains inaccuracies, the quality of generated text, and the quality of references (Methods). It also contains the overall quality of correction by taking all 13 evaluation criteria into account (Methods).
@@
Our primary finding is that the overall quality of Muse-generated corrections is higher than corrections by GPT-4 and even high-helpfulness corrections by laypeople (Fig. 2a). The overall quality of Muse-generated corrections has an average score of 8.1 out of 10, 29% higher than laypeople’s high-helpfulness corrections (mean: 6.3; $p=3\times 10^{-48}$, by Mann-Whitney U test unless otherwise specified; $N=464$), 37% higher than GPT-4-generated corrections (mean: 5.9; $p=4\times 10^{-42}$; $N=464$), and 56% significantly higher than laypeople’s average-helpfulness corrections (mean: 5.2; $p=5\times 10^{-81}$; $N=462$). Despite statistical insignificance between the overall quality of laypeople’s high-helpfulness corrections and GPT-4-generated corrections ($p=0.4$; $N=464$), the overall quality of GPT-4-generated corrections has the highest variability, and GPT-4 generates more corrections with extremely low quality. The standard deviation of the overall quality of GPT-4-generated corrections is 2.7, vs only 2.0 for Muse and laypeople. 10% of GPT-4’s generated corrections have a quality score of 0 (lowest) or 1 out of 10, whereas this proportion is 5% for laypeople’s average-helpfulness corrections, 3% for laypeople’s high-helpfulness corrections, and 2% for Muse-generated corrections. Note that laypeople’s corrections were created on average 14 hours after the tweet was posted on social media. Here, Muse only retrieved web pages published before the tweet was posted (Methods).
@@
Examining specific components of correction quality, results show that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness corrections in identifying and explaining inaccuracies (Fig. 2b-f). Experts assessed that Muse-generated corrections more explicitly identify and explain where and why the tweets are inaccurate than GPT-4’s and laypeople’s high-helpfulness corrections (Fig. 2b). 89% of Muse’s generated corrections explicitly identify and explain inaccuracies, 16% more than GPT-4-generated corrections, 29% more than laypeople’s high-helpfulness corrections, and 43% more than laypeople’s average-helpfulness corrections (Fig. 2b). As for identifying where the tweets are inaccurate, we found that Muse more comprehensively identifies the tweets’ inaccuracies with fewer mistakes (i.e., claiming where the tweets should be inaccurate as accurate or vice versa) than GPT-4 and laypeople who produce even high-helpfulness corrections (Fig. 2c-d). 91% of Muse’s generated corrections have at least one accurate identification without any mistake, 11% more than laypeople’s high-helpfulness corrections, 19% more than GPT-4-generated corrections, and 26% more than laypeople’s average-helpfulness corrections (Fig. 2c). Muse has 61% of generated corrections accurately identifying all the tweets’ inaccuracies, vs GPT-4 has 38%, laypeople who produce high-helpfulness corrections have 26%, and laypeople who produce average-helpfulness correction have 17% only (Fig. 2d). Furthermore, Muse explains why the tweets are inaccurate more precisely and informatively than GPT-4 and laypeople who produce even high-helpfulness corrections (Fig. 2e-f). 70% of corrections by Muse have fully accurate explanations, vs 55% for laypeople’s high-helpfulness corrections, 47% by GPT-4, and 37% for laypeople’s average-helpfulness corrections only (Fig. 2e). Meanwhile, the average informativeness score of Muse-generated corrections is 8.4, 18% higher than GPT-4-generated corrections, 29% higher than laypeople’s high-helpfulness corrections, and 50% higher than laypeople’s average-helpfulness corrections (Fig. 2f).
@@
Results also demonstrate that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness corrections in the quality of generated text (Fig. 2g-k). Muse, which can augment GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals, exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text in correcting misinformation compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$). Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=464$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness corrections by laypeople and additionally less toxic than the text of average-helpfulness corrections by laypeople ($p=4\times 10^{-12}$; $N=462$). In particular, Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness corrections by laypeople, and 43% higher than the text of average-helpfulness corrections by laypeople (Fig. 2g). 74% of Muse-generated text is completely factual, vs 59% for the text of even high-helpfulness corrections by laypeople and 45% for GPT-4-generated text (Fig. 2h). Muse has 99% generated text without any mistake in the use of English, vs laypeople who produce even high-helpfulness corrections have 89% such text (Fig. 2i). Around 91% of Muse-generated text is highly coherent and logical, vs 76% and 61% for the text of high- and average-helpfulness corrections by laypeople (Fig. 2j). Around 99% of Muse-generated text is not biased, impolite, and provoking, 12% higher than that of average-helpfulness corrections by laypeople (Fig. 2k).
@@
Additionally, results reveal that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness corrections in the quality of references (Fig. 2l-n). First, GPT-4 hallucinates references frequently. 49% of its links result in “page-not-found” errors (Fig. 2l), and only 76% of reachable links support or are relevant to the generated text (Fig. 2m). Muse significantly reduces such hallucinations with nearly 100% links being reachable (Fig. 2l) and 96% reachable links being relevant to the generated text (Fig. 2m). Meanwhile, Muse’s references are more credible than the references offered in even high-helpfulness corrections by laypeople ($p=4\times 10^{-11}$; $N=744$; Fig. 2n).
@@
Furthermore, we observed that the quality of Muse-generated corrections to text-only and multimodal (textual and visual) misinformation is higher than corrections by GPT-4 and even high-helpfulness corrections by laypeople (Fig. 3a; Supplementary Fig. S5-S6). The quality of Muse-generated corrections to text-only tweets is 29% higher than GPT-4-generated corrections ($p=6\times 10^{-28}$; $N=310$), 33% higher than laypeople’s high-helpfulness corrections ($p=10^{-44}$; $N=310$), and 65% higher than laypeople’s average-helpfulness corrections ($p=10^{-67}$; $N=310$). The quality of Muse-generated corrections to multimodal tweets is generally 21% higher than laypeople’s high-helpfulness corrections ($p=3\times 10^{-8}$; $N=154$), 39% higher than laypeople’s average-helpfulness corrections ($p=10^{-16}$; $N=154$), and 56% higher than GPT-4 ($p=6\times 10^{-18}$; $N=152$).

### 06 Apr 2024 v2

@@ Abstract
Real-world misinformation can be partially correct and even factual but misleading. It undermines public trust in science and democracy, particularly on social media, where it can spread rapidly. High-quality and timely correction of misinformation that identifies and explains its (in)accuracies has been shown to effectively reduce false beliefs. Despite the wide acceptance of manual correction, this is difficult promptly correct newly created misinformation and to scale this approach, a concern as technologies like large language models (LLMs) make misinformation easier to produce. LLMs also have versatile capabilities that could accelerate misinformation correction—however, they struggle due to a lack of recent information, a tendency to produce false content, and limitations in addressing multimodal information. We propose Muse, an LLM augmented with access to and credibility evaluation of up-to-date information. By retrieving evidence as refutation or supporting context, Muse identifies and explains (in)accuracies in a piece of content—not presupposed to be misinformation—with references. It also describes images and conducts multimodal searches to verify and correct multimodal content. Fact-checking experts evaluate responses to social media content that are not presupposed to be (non-)misinformation but broadly include incorrect, partially correct, and correct posts, that may or may not be misleading. We propose and evaluate 13 dimensions of misinformation correction quality, ranging from the accuracy of identifications and factuality of explanations to the relevance and credibility of references. The results demonstrate Muse’s ability to promptly write high-quality responses to potential misinformation on social media—overall, Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29%. This work reveals LLMs’ potential to help combat real-world misinformation effectively and efficiently.
@@ Introduction
Misinformation, which broadly includes partially incorrect and factual but misleading content[1, 2], has far-reaching and detrimental effects on individuals and society[3, 1, 2, 4, 5]. It erodes public trust in government, decreases civil engagement in elections, and has been viewed as a threat to democracy[6, 2, 7]. Evidence has demonstrated that election misinformation helped fuel the January 6th attack on the U.S. Capitol, where five people died and more than 100 police officers were injured[8]. Misinformation also drastically increases during outbreaks and disasters, as seen with the “infodemic” during the COVID-19 pandemic, which significantly increased vaccine hesitancy[9, 10]. The result is tragic, where COVID-19 vaccines could have prevented at least 232,000 of COVID-19-associated deaths between May 2021 and September 2022 in the U.S. alone[11]. Concerns over misinformation on social media have been particularly significant[12, 13, 14], as the social media context interferes with truth discernment, where users post content without professional moderation and often consume news in a hasty and distracted way[15, 16].
Fortunately, high-quality and timely correction of misinformation, which identifies what part(s) of its content is or is not accurate and explains why that part of the content is (in)accurate with references, has been shown to effectively reduce the spread of misinformation and false beliefs[17, 18, 4, 19]. While domain experts (e.g., on FactCheck.org) and groups of laypeople (e.g., on X Community Notes, formerly Twitter Birdwatch) have played pivotal roles in correcting misinformation[19, 20, 17], keeping pace with massive social media posts is impossible. As Brandolini’s law indicates[21], correcting misinformation is laborious—often requiring one to search for related and trusted articles and write justifications—whereas creating and spreading misinformation is easy. As a consequence, a 88% of suspicious content on X did not receive any response, and 93% did not receive a high-quality response within the first hour (according to X/Twitter Community Notes as of February 2023; Supplementary Fig. S1). Even high-quality responses suffer from limited effectiveness, when they are created after rather than before initial bursts of attention to misinformation content[4, 14]. The absence of sufficient coverage also markedly diminishes the impact of correction and, as the implied truth effect suggests, may even increase the perceived accuracy of misinformation that escapes correction[22].
@@
While generative AI models such as LLMs (e.g., OpenAI’s GPT series) raise concerns that they facilitate creating misinformation, they also potentially make scaling up and accelerating misinformation’s correction possible. Recent LLMs have exhibited proficiency in generating fluent and coherent text, laying a foundation for producing explanations that the public can understand. Indeed, LLMs have revolutionized the field of AI and presented remarkable capabilities across domains and tasks[23, 24, 25]. However, accurate and trustworthy misinformation correction is inseparable from accessing up-to-date and reliable information, providing accurate references to back up claims, and addressing textual and visual information, all areas where existing LLMs fall short[26]. For instance, GPT-4[27] (as of March 2023) and MisinfoCorrect[28] lack access to timely updated knowledge and are thereby ill-equipped to combat misinformation on emerging topics. They either do not provide or “hallucinate” references, which can be fabricated or irrelevant[26, 29]. A growing body of literature has focused on retrieval-augmented LLMs, which can retrieve up-to-date information from Wikipedia or the entire Internet[30, 31, 32, 33, 34, 35, 36, 37]. However, their retrieval does not explicitly consider the factuality and bias of retrieved sources, posing risks of generating falsehood and backfiring (i.e., reinforcing rather than reducing false beliefs)[31, 38, 39]. These LLMs, as well as MisinfoCorrect[28], also struggle with counteracting multimodal misinformation due to their nonacceptance of visual inputs. Finally, understanding the quality of a correction necessitates a comprehensive evaluation of fact-checking experts due to the complexity of real-world misinformation[26]. However, prior evaluation of LLMs was conducted by non-experts on limited aspects of their generated responses (e.g., the general quality of responses[28])[37], which is insufficient to comprehensively understand their performance in identifying and explaining (in)accuracies, generated text, and references.
@@
In this article, we propose Muse, a scalable approach for multimodal misinformation correction. Muse makes use of an LLM and augments it with the ability to handle images, access timely and credible knowledge on the web, retrieve evidence that refutes or contextualizes the given content that may or may not be misinformation, and generate clear explanations with accurate and trustworthy references. Fact-checking experts comprehensively evaluate Muse-generated responses to real social media posts that potentially are misinformation, and compare them to baselines including GPT-4 and high-quality responses based on the collective efforts of laypeople. Our assessment measures the overall quality of a response, specifically defined as the explicitness, accuracy, comprehensiveness, and informativeness when it identifies and explains (in)accuracies, the relevance, factuality, fluency, coherence, and toxicity of generated text, and the reachability, relevance, and credibility of references. We find that Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29% in effectively and promptly responding to potential misinformation. Results demonstrate Muse’s advance when the content is textual with or without images across a broad range of domains, including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine.
@@
Muse is designed to automatically respond to content that potentially is misinformation. In other words, the content might be inaccurate, partially accurate, or even factually accurate but misleading, all of which are misinformation, or fully accurate as non-misinformation. The content can contain text with or without visual information. The response should identify what part(s) of the content is (in)accurate, explain why that part of the content is (in)accurate, and provide links as references. The pipeline of Muse is illustrated in Fig. 1.
@@
We start by introducing the details of Muse with a piece of text-only potential misinformation as the input. First, Muse generates queries based on an LLM from the potential misinformation (Fig. 1b; Methods). Each query acts as the input of a web search engine to access timely updated web content and obtain a list of web links directly relevant to the query (Fig. 1b; Methods). After scraping the content from these web links, Muse calculates their direct relevance to the potential misinformation and removes irrelevant web pages (Fig. 1b; Methods). Then, Muse determines the credibility of web pages by looking up their publishers’ factuality and bias ratings and selects pages with high factuality and minimal bias (Fig. 1c; Methods). Next, Muse leverages an LLM to extract text from each of the web pages as evidence. Such evidence can refute the potential misinformation, typically happening when it is misinformation with false claims, or provide additional context, which can demonstrate that the potential misinformation is accurate or part(s) of its claims are accurate (Fig. 1d; Methods). Finally, Muse generates a response to the potential misinformation by providing an LLM with the extracted pieces of evidence and their source web links (Fig. 1d; Methods).
@@
Note that content, especially content posted on social media, often contains extraneous information that does not need verification and is irrelevant to correction, including unverifiable opinions or emojis, such as the textual content of the false post in Fig. 1. Therefore, generating queries instead of simply using the post content improves web searches as a way of denoising the post content; see example queries in Fig. 1b. Meanwhile, generating multiple queries helps decompose a post, which may have multiple claims that each needs verification or correction, whereas generating one query may overlook some of the claims and hence lead to not comprehensive identifications and explanations of (in)accuracies (Supplementary Fig. S2). Another concern may arise from filtering retrieved web pages based on how relevant the page content is to the potential misinformation: theoretically, including all retrieved web pages increases the amount of extracted evidence, which may not hurt and perhaps even benefit correction. However, the increase in selected web pages drastically elevates the expense of Muse (Methods). We also observe that retrieved web pages with relatively low relevance can increase the prevalence of hallucinations when generating responses (Supplementary Fig. S3). Moreover, as illustrated in Fig. 1c, Muse filters and ranks the selected web pages by their publishers’ factuality and bias. It starts extracting evidence from pages with the highest factuality and least bias and then continues down the ranking, stopping when it has obtained sufficient refutations (i.e., at least two web pages were found to refute the misinformation) or gone through all the credible pages.
@@
When the input is a piece of multimodal (textual and visual) content, Muse first generates textual description of each image (Fig. 1a) so that the content can be handled by any LLM on downstream tasks, including query generation (Fig. 1b), evidence extraction (Fig. 1d), and response generation (Fig. 1d). Specifically, Muse augments image captioning models developed to describe an image in natural language with recognizing celebrities and optical characters based on an LLM (Fig. 1a; Methods). Compared to existing image captioning models that capture global features of images, Muse produces more informative descriptions with additional features, including identification of celebrities and embedded text—information crucial for making accurate verification and corrections[40]. For example, even a state-of-the-art image captioning model[41] may describe the visual misinformation in Fig. 1 as simply “list of banned books in Florida.” The description overlooks the listed titles of books that are essential for the visual content’s verification and correction (see more examples in Supplementary Fig. S4). In addition, Muse conducts multimodal search on the web and computes multimodal relevance to filter out irrelevant web pages (Fig. 1b; Methods).
@@
Muse’s evaluation was based on X Community Notes data. Community Notes empowers people on X, often laypeople, to collaboratively fact-check tweets, which has shown to reduce the spread of misinformation[17]. Every laypeople’s free-response fact-check is associated with a helpfulness score by aggregating the assessments of people with diverse backgrounds, e.g., different political ideologies. A response with a sufficiently high helpfulness score is then displayed on the corresponding tweet and publicly visible[17] (Supplementary Fig. S6). We included the tweets from Community Notes with at least two responses (#=247); one has a high helpfulness score and the other has an average helpfulness score (as of February 2023; Methods). Though we do not presuppose the accuracy of the tweets in Muse’s design and evaluation, we found that more than half of the tweets are not fully (in)accurate or misleading but frequently presented in a way that combines accurate claims and inaccurate or misleading claims (Methods). We further generated responses to these tweets based on Muse and GPT-4 (as of June 2023) (Discussion; Methods). Experts in fact-checking and journalism evaluated the quality of responses by various approaches to the same tweet (Methods); they were blinded to which approach had generated each response. The evaluation contains 13 specific criteria, covering how well a response identifies and explains (in)accuracies, the quality of generated text, and the quality of references (Methods). It also contains the overall quality of a response by taking all 13 evaluation criteria into account (Methods).
@@
Our primary finding is that the overall quality of Muse-generated responses is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig. 2a). The overall quality of Muse-generated responses has an average score of 8.1 out of 10, 29% higher than laypeople’s high-helpfulness responses (mean: 6.3; $p=3\times 10^{-48}$, by Mann-Whitney U test unless otherwise specified; $N=464$), 37% higher than GPT-4-generated responses (mean: 5.9; $p=4\times 10^{-42}$; $N=464$), and 56% significantly higher than laypeople’s average-helpfulness responses (mean: 5.2; $p=5\times 10^{-81}$; $N=462$). Despite statistical insignificance between the overall quality of laypeople’s high-helpfulness responses and GPT-4-generated responses ($p=0.4$; $N=464$), the overall quality of GPT-4-generated responses has the highest variability, and GPT-4 generates more responses with extremely low quality. The standard deviation of the overall quality of GPT-4-generated responses is 2.7, vs only 2.0 for Muse and laypeople. 10% of GPT-4’s generated responses have a quality score of 0 (lowest) or 1 out of 10, whereas this proportion is 5% for laypeople’s average-helpfulness responses, 3% for laypeople’s high-helpfulness responses, and 2% for Muse-generated responses. Note that laypeople’s responses were created on average 14 hours after the tweet was posted on social media. Here, Muse only retrieved web pages published before the tweet was posted (Methods).
@@
Examining specific components of response quality, results show that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in identifying and explaining (in)accuracies (Fig. 2b-f). Experts assessed that Muse-generated responses more explicitly identify and explain where and why a tweet is (in)accurate than GPT-4’s and laypeople’s high-helpfulness responses (Fig. 2b). 89% of Muse’s generated responses explicitly identify and explain (in)accuracies, 16% more than GPT-4-generated responses, 29% more than laypeople’s high-helpfulness responses, and 43% more than laypeople’s average-helpfulness responses (Fig. 2b). As for identifying where a tweet is (in)accurate, we found that Muse more comprehensively identifies a tweet’s (in)accuracies with fewer mistakes—here, mistakes indicate falsely claiming where a tweet should be inaccurate as accurate or where a tweet should be accurate as inaccurate—than GPT-4 and laypeople who produce even high-helpfulness responses (Fig. 2c-d). 91% of Muse’s generated responses have at least one correct identification without any mistake, 11% more than laypeople’s high-helpfulness responses, 19% more than GPT-4-generated responses, and 26% more than laypeople’s average-helpfulness responses (Fig. 2c). Muse has 61% of generated responses accurately identifying all the (in)accuracies in a tweet, vs GPT-4 has 38%, laypeople who produce high-helpfulness responses have 26%, and laypeople who produce average-helpfulness responses have 17% only (Fig. 2d). Furthermore, Muse explains (in)accuracies more precisely and informatively than GPT-4 and laypeople who produce even high-helpfulness responses (Fig. 2e-f). 70% of responses by Muse have fully accurate explanations, vs 55% for laypeople’s high-helpfulness responses, 47% by GPT-4, and 37% for laypeople’s average-helpfulness responses only (Fig. 2e). Meanwhile, the average informativeness score of Muse-generated responses is 8.4, 18% higher than GPT-4-generated responses, 29% higher than laypeople’s high-helpfulness responses, and 50% higher than laypeople’s average-helpfulness responses (Fig. 2f).
@@
Results also demonstrate that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in the quality of generated text (Fig. 2g-k). Muse, when it augments GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals (Methods), exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$). Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=464$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness responses by laypeople and additionally less toxic than the text of average-helpfulness responses by laypeople ($p=4\times 10^{-12}$; $N=462$). In particular, Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness responses by laypeople, and 43% higher than the text of average-helpfulness responses by laypeople (Fig. 2g). 74% of Muse-generated text is completely factual, vs 59% for the text of even high-helpfulness responses by laypeople and 45% for GPT-4-generated text (Fig. 2h). Almost all of Muse-generated text does not have any mistake in the use of English (Fig. 2i) and is not biased, impolite, and provoking (Fig. 2k), and 91% is highly coherent and logical (vs 76% and 61% for laypeople, Fig. 2j).
@@
Additionally, results reveal that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in the quality of references (Fig. 2l-n). First, GPT-4 hallucinates references frequently. 49% of its links result in “page-not-found” errors (Fig. 2l), and only 76% of reachable links are relevant to the generated text (Fig. 2m). Muse significantly reduces such hallucinations with nearly 100% links being reachable (Fig. 2l) and 96% reachable links being relevant to the generated text (Fig. 2m). Meanwhile, Muse’s references are more credible than the references offered in even high-helpfulness responses by laypeople ($p=4\times 10^{-11}$; $N=744$; Fig. 2n).
@@
Furthermore, we observed that the quality of Muse-generated responses to textual and multimodal (textual and visual) content that potentially is misinformation is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig. 3a; Supplementary Fig. S7-S8). The quality of Muse-generated responses to text-based tweets is 29% higher than GPT-4-generated responses ($p=6\times 10^{-28}$; $N=310$), 33% higher than laypeople’s high-helpfulness responses ($p=10^{-44}$; $N=310$), and 65% higher than laypeople’s average-helpfulness responses ($p=10^{-67}$; $N=310$). The quality of Muse-generated responses to multimodal tweets is generally 21% higher than laypeople’s high-helpfulness responses ($p=3\times 10^{-8}$; $N=154$), 39% higher than laypeople’s average-helpfulness responses ($p=10^{-16}$; $N=154$), and 56% higher than GPT-4 ($p=6\times 10^{-18}$; $N=152$).